OpenStudy (anonymous):

For each function below, determine A. f (x-1), B. f (x+a)-f (x), and C. f (x^2).

OpenStudy (anonymous):

f (x)= 3x+2

OpenStudy (experimentx):

f(x-1) = 3(x-1)+2 = 3x-1 f(x+a) - f(x) = 3(x+a)+2 - 3(x)-2 = .. f(x^2) = 3(x^2)+2

OpenStudy (anonymous):

Now do i just solve it?

OpenStudy (experimentx):

find the second one

OpenStudy (anonymous):

f(x+a)... Is that the one your talking bout?